# Closed-Loop Neural Operator-Based Observer of Traffic Density

### v1

@@ Abstract
We consider the problem of traffic density estimation with sparse measurements from stationary roadside sensors. Our approach uses Fourier neural operators to learn macroscopic traffic flow dynamics from high-fidelity microscopic-level simulations. During inference, the operator functions as an open-loop predictor of traffic evolution. To close the loop, we couple the open-loop operator with a correction operator that combines the predicted density with sparse measurements from the sensors. Simulations with the SUMO software indicate that, compared to open-loop observers, the proposed closed-loop observer exhibit classical closed-loop properties such as robustness to noise and ultimate boundedness of the error. This shows the advantages of combining learned physics with real-time corrections, and opens avenues for accurate, efficient, and interpretable data-driven observers.
@@ I Introduction
Freeway congestion is a major issue in metropolitan areas, leading to increased travel times and excessive fuel consumption [1]. To address this, a core component of Intelligent Transportation Systems is freeway traffic control, which operates at the macroscopic level (ramp management and variable speed limits) or at the vehicle level (intelligent vehicle-based control).
Traffic control relies on measurements of state variables such as traffic density, which can be collected from stationary sensors (inductive loops) or mobile sensors (connected vehicles). However, these measurements are often sparse and noisy. To complete and denoise the state observation, estimation approaches combine data and prior knowledge of traffic dynamics [2].
@@
Model-driven methods use physical models with calibrated parameters. Traffic state estimation then boils down to solving boundary-value problems of partial differential equations (PDEs) using numerical methods such as the Godunov scheme [3] or switching mode models [4]. To manage measurement noise and disturbance, it is common to use Kalman Filter approaches [5]. While these methods require little data and offer high interpretability, their accuracy is highly dependent on model choice and calibration [2].
@@
Data-driven methods, on the other hand, learns a system model directly from historical data, thereby alleviating explicit PDE assumptions and requirements of well-posed boundary-value problems [2]. Classical methods include ARIMA variants, Gaussian process (GP) models, and state-space Markov models. The rise of deep learning for nonlinear function approximation has enabled more expressive models of traffic flow prediction [6]. In particular, Convolutional Neural Networks (CNNs) and Recurrent Neural Networks (RNNs) are often used as building blocks to capture spatiotemporal correlation. While deep learning approaches are expressive and require few assumptions, they are limited by a high data demand and computational complexity, and low interpretability. Thus, lightweight real-time prediction with explainable models remains an open challenge [7].
@@
In pursuit of data-efficient and interpretable models, hybrid models combine data-driven methods with prior physical knowledge. For example, [8] reconstructs the traffic density state using physics-informed neural networks (PINNs) trained on offline data from sparse mobile sensors while regularizing with a continuous traffic PDE model. In [9], traffic models are brought into a GP-based architecture via latent force models and regularized kernel learning. However, these methods are restricted by explicit model assumptions on global traffic behavior. Furthermore, online estimation can be computationally demanding, since it requires retraining the network (PINNs) or performing matrix inversion with cubic complexity relative to the size of the dataset (GPs).
@@
We identify the need for a data-driven observer of traffic density that encodes a learned governing model in an interpretable manner, offers lightweight real-time prediction, and integrates online measurements. For finite-dimensional nonlinear systems, there are learning-based solutions in the form of closed-loop data-driven observers where training is done offline and low-cost inference is performed online [10]. The natural extension to PDE systems would be to consider Fourier neural operators (FNOs) [11] which has emerged as a method to learn operators between infinite-dimensional function spaces from finite-dimensional data.
@@
Notably, [12] successfully applied FNOs to learn a PDE solution operator for initial-boundary value problems and inverse problems of a first-order macroscopic model of traffic density flow. With this method, inference with new input data is significantly cheaper than with PINNs, since it is performed by a single forward pass through the operator network. However, the solver in this work is designed to generate either an offline reconstruction or open-loop prediction, without systematic integration of online data.
@@
Motivated by this, we introduce a data-driven, closed-loop neural observer of traffic density flow from sparse measurements. Our contributions are leveraging FNOs to learn a prediction operator of traffic density flow from high-fidelity data, and integrating Luenberger observer theory to get a robust density estimate using online measurements. We statistically evaluate the performance of the observer and compare it to open-loop variants.
@@
Notation: Let $\mathcal{H}^{d}_{D}\triangleq H^{2}_{\text{per}}(D;\mathbb{R}^{d})$ denote the Sobolev space of order two containing periodic functions with domain $D$ and codomain dimension $d$. Further, we denote an operator with $\mathcal{G}$ or $\mathcal{N}$. Finally, a Gaussian Process (GP) is denoted $\mathcal{GP}$, and the posterior process associated with data $\mathcal{D}$ is denoted $\mathcal{GP}_{\mathcal{D}}$. We assume that the mean and covariance functions are chosen such that sample paths belong to $\mathcal{H}^{d}_{D}$ almost surely.
@@ II-A1 Macroscopic scale
A fundamental continuous first-order macroscopic model for traffic density flow is given by the Lighthill–Whitham–Richards framework. It expresses conservation of vehicles through the one-dimensional continuity equation where the flux $Q\in C^{2}([0,1])$ is uniquely determined by the local density [13]:
@@
In general, classical solutions to the initial-value problem of the hyperbolic PDE (1) where $Q$ is smooth may develop discontinuities in finite time even when the initial data $\rho_{0}$ is smooth [14]. The weak solutions, on the other hand, are non-unique but can be filtered for physical relevancy using an entropy condition associated with the conservation law (1). In fact, whenever $Q$ is smooth and $\rho_{0}$ is well-behaved, there exist a unique admissible (entropic) solution $\rho(\cdot,t)$ that is continuous in time and locally well-behaved [14, Theorem 6.2.1, 6.2.2]. In this view, traffic density estimation is a well-defined problem. However, a key challenge for macroscopic models is to properly model the flow function $Q$.
@@ II-A2 Microscopic scale
To generate representations of higher fidelity without direct modeling of $Q$, traffic flow can instead be simulated using higher-order models at the vehicle level[15]. The open-source SUMO framework [16] integrates road network data, infrastructure like traffic lights, and stochastic demand to produce realistic vehicle flows. In particular, we will consider ring-road dynamics as implemented by [17]: first, the initialization process gradually introduces vehicles to the road to create the initial state; then, the higher-order microscopic dynamics are run in loop. To convert the discrete distribution of vehicles to a continuous density function, a grid of vehicle-populated cells is convolved with a Gaussian filter [18]. In this way, microscopic models generate high-fidelity traffic representations with few restrictive assumptions. However, a key challenge is the lack of a compact representation of the dynamics that effectively integrates with macroscopic-level data and control strategies.
@@ II-B Fourier Neural Operators (FNOs)
To address this challenge, neural operators provide a framework for learning the governing dynamics on the macroscopic level. Generally, neural operators learn a mapping between infinite-dimensional function spaces, that is spaces containing functions with continuous domains, using finite-dimensional data. A special case is when the operator represents a forward solver for a hyperbolic PDE, such that the input function is the initial state and the output function is the solution at a final time $T$.
We introduce neural operators in the context of the traffic density flow arising from the higher-order, microscopic dynamics in SUMO simulations on a ring road $\Omega=[0,L]$. Consider pairs of functions
$(\zeta_{0},\rho_{T})\in\mathcal{H}^{n}_{\Omega}\times\mathcal{H}^{1}_{\Omega},$
where $\zeta_{0}=[\rho(\cdot,0),\rho(\cdot,-\Delta t),\rho(\cdot,-2\Delta t),...]$ represents the discrete $n^{\text{th}}$-order initial state of step length $\Delta t$, and $\rho_{T}=\rho(\cdot,T)$ represents the resulting traffic density at time $T$. Assume there exists an operator $\mathcal{G}$ such that
$\mathcal{G}:\begin{array}[t]{rcl}\mathcal{H}^{n}_{\Omega}&\to&\mathcal{H}^{1}_{\Omega},\\ \zeta_{0}&\mapsto&\rho_{T},\end{array}$
as can be motivated by the existence of a solution in the first-order case (1). Given observations of such pairs $\{(\zeta_{0}^{i},\rho_{T}^{i})\}_{i=1}^{N}$ resulting from simulations initialized by $\zeta_{0}^{i}\sim\mu_{\text{SUMO}}$, the goal is to learn an approximate solution operator $\mathcal{G}_{\theta}\approx\mathcal{G}$ where $\mathcal{G}_{\theta}$ is parametrized by finite-dimensional vector $\theta$.
The approximate solution operator $\mathcal{G}_{\theta}:\mathcal{H}^{n}_{\Omega}\to\mathcal{H}^{1}_{\Omega}$ can be modeled by a neural operator as defined in [19]:
$\begin{split}v_{1}(x)=\mathcal{P}[\zeta_{0}](x),&\\ v_{M}(x)=\mathcal{L}_{M}\circ...\circ\mathcal{L}_{1}[v_{1}](x),&\\ \rho_{T}(x)=\mathcal{Q}[v_{M}](x),\quad\forall x\in\Omega,&\end{split}$ (2)
where $\mathcal{P}$ is defined by a local transformation $P:\mathbb{R}^{n}~{}\times~{}\Omega\to\mathbb{R}^{d_{v}}$, $\mathcal{P}[\zeta_{0}](x)\triangleq P(\zeta_{0}(x),x)$ acting on function inputs and outputs directly to lift the input state to higher-dimensional space of size $v$. Similarly, $\mathcal{Q}$ is defined by a local transformation $Q:\mathbb{R}^{d_{v}}\times\Omega\to\mathbb{R}$, $\mathcal{Q}[v](x)\triangleq Q(v(x),x)$ projecting back to the physical space. Both are typically shallow neural networks. The intermediate spectral convolution layers $\mathcal{L}_{l}:\mathcal{V}\to\mathcal{V}$, where $\mathcal{V}=\mathcal{V}(\Omega;\mathbb{R}^{d_{v}})$, are structured as
@@
where $\sigma$ is a local nonlinear activation function, and $W_{l}$ is a local pointwise affine transformation. The key component for the infinite-dimensional setting is the non-local bounded linear operator $\mathcal{K}_{\phi_{l}}\in\mathcal{L}(\mathcal{V},\mathcal{V})$, which acts in the function space.
FNOs, proposed by [11], select $\mathcal{K}_{\phi_{l}}$ as a kernel integral operator such that
@@
where $\kappa_{\phi_{l}}:\bar{\Omega}\to\mathbb{R}^{d_{v}\times d_{v}}$ is a periodic function. The choice (3) allows for efficient computation in the Fourier domain via the equivalent expression
$\mathcal{K}_{\phi_{l}}[v](x)=\mathcal{F}^{-1}\big{(}\mathcal{F}(\kappa_{\phi_{l}})(k)\cdot\mathcal{F}(v)(k)\big{)}(x)$ (4)
where $\mathcal{F}$ is the Fourier transform and $k\in\mathbb{Z}$. The kernel can then be parametrized directly in the Fourier space as $R_{\phi_{l}}\triangleq\mathcal{F}(\kappa_{\phi_{l}})$. We will work with a uniform discretization of $\Omega$, thus $\mathcal{F}$ may be replaced by Fast Fourier Transform $\hat{\mathcal{F}}$.
Learning the parameters in the frequency domain, as opposed to the physical domain, enables a finite-dimensional representation of the operator while maintaining its ability to map between infinite-dimensional function spaces. The finite-dimensional representation is obtained by truncating the Fourier expansions to $k_{max}$ frequency modes, such that ^$\hat{\mathcal{F}}(v)\in\mathbb{C}^{k_{max}\times d_{v}}$ and $R_{\phi_{l}}\in\mathbb{C}^{k_{max}\times d_{v}\times d_{v}}$. Still, (4) is defined on a continuous domain, since ^$\hat{\mathcal{F}}^{-1}$ can project on the basis $e^{2\pi i\langle\cdot,k\rangle}$ for any $x\in\Omega$. In practice, this allows us to learn mappings between function spaces with continuous domains even though the data consists of function evaluations on discretized grids. The resulting architecture $\mathcal{G}_{\theta}$ is fast and achieves state-of-the art results for PDE problems [11].
@@
Let $\mathcal{G}:\mathcal{H}^{n}_{\Omega}\to\mathcal{H}^{1}_{\Omega}$ be a continuous operator, and $K\subset\mathcal{H}^{n}_{\Omega}$ a compact subset. Let $\mathcal{G}_{\theta}:\mathcal{H}^{n}_{\Omega}\to\mathcal{H}^{1}_{\Omega}$ be defined as in (2) with the coefficients $\theta\triangleq\{\theta_{P},\theta_{Q},\theta_{W_{l}},R_{\phi_{l}}\}$. Then, $\forall\epsilon>0$ $\exists\theta$ such that $\mathcal{G}_{\theta}$, continuous as an operator, satisfies
[…]

### v2

@@ Abstract
We consider the problem of traffic density estimation with sparse measurements from stationary roadside sensors. Our approach uses Fourier neural operators to learn macroscopic traffic flow dynamics from high-fidelity data. During inference, the operator functions as an open-loop predictor of traffic evolution. To close the loop, we couple the open-loop operator with a correction operator that combines the predicted density with sparse measurements from the sensors. Simulations with the SUMO software indicate that, compared to open-loop observers, the proposed closed-loop observer exhibits classical closed-loop properties such as robustness to noise and ultimate boundedness of the error. This shows the advantages of combining learned physics with real-time corrections, and opens avenues for accurate, efficient, and interpretable data-driven observers.
@@ I Introduction
Freeway congestion is a major issue in metropolitan areas, leading to increased travel times and excessive fuel consumption [1]. To address this, a core component of Intelligent Transportation Systems is freeway traffic control, which operates at the macroscopic level (ramp management and variable speed limits) or at the vehicle level (intelligent vehicle-based control). Traffic control relies on measurements of state variables such as traffic density, which can be collected from stationary sensors (inductive loops) or mobile sensors (connected vehicles). However, these measurements are often sparse and noisy. To complete and denoise the state observation, estimation approaches combine data and prior knowledge of traffic dynamics [2].
@@
Model-based methods use physical models with calibrated parameters. Traffic state estimation then typically boils down to solving initial-boundary value problems (IBVPs) of partial differential equations (PDEs) with numerical methods such as the Godunov scheme [3, 4] or switching mode models [5]. To manage measurement noise and disturbance, it is common to use Kalman or particle filter-based approaches [6, 7]. For example, [8, 9] reformulate the continuity equation into a linear parameter-varying (LPV) system, where real-time flow measurements are treated as parameters. While this class of methods typically requires little data and offers high interpretability, the accuracy is highly dependent on model assumptions and calibration [2].
@@
Data-driven methods, on the other hand, learn a system model directly from historical data, thereby relaxing explicit PDE assumptions and requirements of well-posedness [2]. Classical methods include ARIMA variants, Gaussian process (GP) models, and state-space Markov models. The rise of deep learning for nonlinear function approximation has enabled more expressive models of traffic flow prediction, with early examples such as the stacked autoencoder [10]. In particular, Convolutional Neural Networks (CNNs) and Recurrent Neural Networks (RNNs) are often used as building blocks to capture spatiotemporal correlation. While deep learning approaches are expressive and require few assumptions, they are limited by a high data demand and computational complexity, and low interpretability. Thus, lightweight real-time prediction with explainable models remains an open challenge [11].
@@
In pursuit of data-efficient and interpretable models, hybrid models combine data-driven methods with prior physical knowledge. For example, [12] reconstructs the traffic density state using physics-informed neural networks (PINNs) trained on offline data from sparse mobile sensors while regularizing with a continuous traffic PDE model. In [13], traffic models are used in physics-regularized GPs (PRGPs), which rely on latent force models and regularized kernel learning. However, these methods are restricted by explicit model assumptions on global traffic behavior. Furthermore, online estimation can be computationally demanding, since it requires retraining the network (PINNs) or performing matrix inversion with cubic complexity in terms of the size of the dataset (PRGPs).
@@
We identify the need for a data-driven observer of traffic density that encodes a learned governing model in an interpretable manner, and offers lightweight real-time prediction with the integration of online measurements. For finite-dimensional nonlinear systems, there are learning-based solutions in the form of closed-loop data-driven observers where training is done offline and low-cost inference is performed online [14]. The natural extension to PDE systems would be to consider Fourier neural operators (FNOs) [15], which have emerged as a method to learn operators between infinite-dimensional function spaces from finite-dimensional data.
Model approach Observation mechanism
Method Model Data Learned closure Online inference Closed-loop Grid-free
Numerical integration of IBVP[3, 4, 5]
Filtered state-space model [6, 7]
Filtered LPV-model with data-driven parameters [8, 9]
Stacked autoencoder [10]; CNN + RNN[11]
PINN [12]
PRGP [13]
KKL observer [14]
FNO [16] ✓& ✗ ✓& ✗ ✓& ✗
FNO-based observer ✓& ✗ ✓& ✗ ✓& ✗
TABLE I: Comparison of related methods for traffic state estimation.
@@
Notably, [16] successfully applied FNOs to learn a PDE solution operator for IBVPs and inverse problems of a first-order macroscopic model of traffic density flow. With this method, inference with new input data is significantly cheaper than with PINNs, since it is performed by a single forward pass through the operator network. However, the solver in this work is designed to generate either an offline reconstruction or open-loop prediction, without closed-loop integration of online data.
@@
Motivated by this, we introduce a data-driven, closed-loop neural observer of traffic density flow from sparse measurements. Our contributions are leveraging FNOs to learn a prediction operator of traffic density flow from high-fidelity data, and integrating Luenberger observer theory to get a robust density estimate using online measurements. We statistically evaluate the performance of the observer and compare it to open-loop variants. The relationship between the proposed observer and existing alternative methods is summarized in Table I.
@@
Notation: Let $\mathcal{H}^{d}_{D}\triangleq H^{2}_{\text{per}}(D;\mathbb{R}^{d})$ denote the Sobolev space of order two containing periodic functions with domain $D$ and codomain dimension $d$. Further, we denote an operator with $\mathcal{G}$ or $\mathcal{N}$. Finally, a GP is denoted $\mathcal{GP}$, and the posterior process associated with data $\mathcal{D}$ is denoted $\mathcal{GP}_{\mathcal{D}}$. We assume that the mean and covariance functions are chosen such that sample paths belong to $\mathcal{H}^{d}_{D}$ almost surely.
@@ II-A1 Macroscopic scale
A fundamental continuous first-order macroscopic model for traffic density flow is given by the Lighthill–Whitham–Richards framework. It expresses conservation of vehicles through the one-dimensional continuity equation, where the flow $Q\in C^{2}([0,1])$ is uniquely determined by the local density [17]:
@@
In general, classical solutions to the initial value problem of the hyperbolic PDE (1) where $Q$ is smooth may develop discontinuities in finite time even when the initial data $\rho_{0}$ is smooth [18]. The weak solutions, on the other hand, are non-unique but can be filtered for physical relevancy using an entropy condition associated with the conservation law (1). In fact, whenever $Q$ is smooth and $\rho_{0}$ is well-behaved, there exists a unique admissible (entropic) solution $\rho(\cdot,t)$ that is continuous in time and locally well-behaved [18, Theorem 6.2.1, 6.2.2]. In this view, traffic density estimation is a well-defined problem. However, a key challenge for macroscopic models is to properly model the flow $Q$.
@@ II-A2 Microscopic scale
To generate representations of higher fidelity without direct modeling of $Q$, traffic flow can instead be simulated using higher-order models at the vehicle level[1]. These models integrate interactions between vehicles and road infrastructure to produce realistic traffic flows. Vehicle interactions are commonly represented by car-following models, in which the behavior relative to the preceding vehicle or platoon is specified. For example, [19] proposed to let cars drive as fast as possible while maintaining a safe velocity to avoid collision. The model can be made increasingly complex by introducing parameters such as driver reaction time, estimation errors, and anticipation [20]. In this way, microscopic models may generate high-fidelity traffic flows with few restrictive assumptions. However, a key challenge is the lack of a compact representation of the dynamics that effectively integrates with macroscopic-level data and control strategies.
@@ II-B Fourier Neural Operators (FNOs)
To address this challenge, neural operators provide a framework for learning the governing dynamics on the macroscopic level. Generally, neural operators learn a mapping between infinite-dimensional function spaces, that is, spaces containing functions with continuous domains, using finite-dimensional data. A special case is when the operator represents a forward solver for a hyperbolic PDE, such that the input function is the initial state and the output function is the solution at a final time $T$.
We introduce neural operators in the context of traffic density flow arising from higher-order dynamics on a ring road of length $L$, denoted $\Omega\triangleq[0,L]$. The closed course induces realistic traffic flows, including stop-and-go waves formed by tiny fluctuations [21, 22]. Consider pairs of functions
$(\mathrm{P}_{0},\rho_{T})\in\mathcal{H}^{n}_{\Omega}\times\mathcal{H}^{1}_{\Omega},$
where $\mathrm{P}_{0}=[\rho(\cdot,0),\rho(\cdot,-\Delta t),\rho(\cdot,-2\Delta t),$...] represents the discrete $n^{\text{th}}$-order initial state of step length $\Delta t$, and $\rho_{T}=\rho(\cdot,T)$ represents the resulting traffic density at time $T$. Assume there exists an operator $\mathcal{G}$ such that
$\mathcal{G}:\begin{array}[t]{rcl}\mathcal{H}^{n}_{\Omega}&\to&\mathcal{H}^{1}_{\Omega},\\ \mathrm{P}_{0}&\mapsto&\rho_{T},\end{array}$
as can be motivated by the existence of a solution in the first-order case (1). Given observations of such pairs $\{(\mathrm{P}_{0}^{i},\rho_{T}^{i})\}_{i=1}^{N}$, where the initial states are generated according to some distribution $\mathrm{P}_{0}^{i}\sim\mu$, the goal is to learn an approximate solution operator $\mathcal{G}_{\theta}\approx\mathcal{G}$ where $\mathcal{G}_{\theta}$ is parametrized by a finite-dimensional vector $\theta$.
The approximate solution operator $\mathcal{G}_{\theta}:\mathcal{H}^{n}_{\Omega}\to\mathcal{H}^{1}_{\Omega}$ can be modeled by a neural operator [15]:
[…]
@@
where $\sigma$ is a local nonlinear activation function, and $W_{l}$ is a local pointwise affine transformation. The key component for the infinite-dimensional setting is the non-local bounded linear operator $\mathcal{K}_{\phi_{l}}\in\mathcal{L}(\mathcal{V},\mathcal{V})$, which acts in function space.
A central aspect of FNOs [15] is that $\mathcal{K}_{\phi_{l}}$ is defined by the convolution
@@
where $\kappa_{\phi_{l}}:\Omega\to\mathbb{R}^{d_{v}\times d_{v}}$ is a periodic, integrable function which has a Fourier series representation [^1] This aligns with our problem setting, as the ring road formulation entails that $\rho(\cdot,t)$ is $L$-periodic on $\Omega$.. By the convolution theorem, (3) admits the frequency domain form
$\mathcal{K}_{\phi_{l}}[v](x)=\mathcal{F}^{-1}\big{(}\mathcal{F}(\kappa_{\phi_{l}})(k)\cdot\mathcal{F}(v)(k)\big{)}(x),$ (4)
where $\mathcal{F}$ computes the Fourier coefficients and $k\in\mathbb{Z}$. Parameterizing the convolution kernel directly in the frequency domain, $R_{\phi_{l}}\triangleq\mathcal{F}(\kappa_{\phi_{l}})$, and truncating the Fourier series to $k_{max}$ frequency modes, $R_{\phi_{l}}\in\mathbb{C}^{k_{max}\times d_{v}\times d_{v}}$ and $\mathcal{F}(v)\in\mathbb{C}^{k_{max}\times d_{v}}$, yields a finite-dimensional operator parametrization. Despite this truncation, the operator (3) remains defined on functions over the continuous domain $\Omega$, since $\mathcal{F}^{-1}$ reconstructs from the basis functions $e^{2\pi i\langle\cdot,k\rangle}$ for any $x\in\Omega$. In practice, input–output data consist of discretized rather than analytic functions, and the Fourier transform $\mathcal{F}$ can be computed using the Fast Fourier Transform $\hat{\mathcal{F}}$. Crucially, truncating in the frequency domain – rather than the physical domain – allows for a parametrization that can be shared across resolutions. In other words, the resulting operator (2)–(4) is discretization convergent, and thus consistent in the infinite-dimensional setting.
@@
Let $\mathcal{G}:\mathcal{H}^{n}_{\Omega}\to\mathcal{H}^{1}_{\Omega}$ be a continuous operator, and $K\subset\mathcal{H}^{n}_{\Omega}$ a compact subset. Let $\mathcal{G}_{\theta}:\mathcal{H}^{n}_{\Omega}\to\mathcal{H}^{1}_{\Omega}$ be defined as in (2)–(4) with the coefficients $\theta\triangleq\{\theta_{P},\theta_{Q},\theta_{W_{l}},R_{\phi_{l}}\}$. Then, $\forall\epsilon>0$ $\exists k_{max}\in\mathbb{N}$ and $\theta\in\Theta$ such that $\mathcal{G}_{\theta}$, continuous as an operator, satisfies
$\sup_{a\in K}\int\limits_{\Omega}\|\mathcal{G}[a](x)-\mathcal{G}_{\theta}[a](x)\|^{2}dx\leq\epsilon.$ (5)
It is further suggested by [23] that, while the approximation error scales super-exponentially with the size of the FNO in the general case, it typically scales polynomially for PDEs of the form (1) under suitable assumptions. The expression (5) serves as the basis for a training loss function, and may be extended to include PINN-type regularization from a predefined PDE model [24].